OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x - cos x _______________________________________________ a. 2 sin x - sin3x - cos x b. 2 sin x cos2x + sin x - 2 sin3x - cos x c. 2 sin3x cos4x + 1 d. 3 sin x cos x - sin3x - cos x

jimthompson5910 (jim_thompson5910):

is that sin^3 or is that sin(3x)

OpenStudy (anonymous):

sin(3x)

jimthompson5910 (jim_thompson5910):

so you can use the identity sin(A+B) = sin(A)cos(B) + cos(A)sin(B)

jimthompson5910 (jim_thompson5910):

sin(3x) = sin(x+2x) sin(3x) = sin(x)cos(2x) + cos(x)sin(2x) sin(3x) = sin(x)cos(2x) + cos(x)(2*sin(x)*cos(x)) ... use identity sin(2A) = 2sin(A)cos(A) keep going

OpenStudy (anonymous):

im a bit stuck

OpenStudy (anonymous):

when i did sin(A+B) = sin(A)cos(B) + cos(A)sin(B), my result was 0

OpenStudy (anonymous):

\[\sin 3x=\sin \left( 2x+x \right)=you can solve\] again sin 2x =2 sin x cos x \[\cos 2x=\cos ^{2}x -\sin ^{2}x\]